# FIG: Forward-Inverse Generation for Low-Resource Domain-specific Event Detection

### v1

@@ B.1 Data Statistics
Table 8: Data Statistics for the various test and unlabeled datasets used in our work.
@@ Appendix C Implementation Details
Here, we provide detailed implementation details for each component and models used in our work. We run most of our experiments on NVIDIA RTX A6000/A100 machines with support for 8 GPUs, while for GPT3.5, we make API calls through OpenAI.
@@ C.1 LLM-based Generation
We provide details about the various hyperparameters for using LLMs in all the components of STAR and FIG. For Llama3-8B-Instruct and Llama3-70B-Instruct, we present the hyperparameters in Table 9; while Table 10 presents the hyperparameters for GPT3.5.
@@ C.2 Few-shot Implementation Details
For the few-shot setting, we can access additional $k$ datapoints per event type to aid better performance. For LLM-based prompting, we simply add these examples in the prompt as in-context examples to help the model do better reasoning/generation. For inverse generation methods (STAR, FIG), we do not add the $k$ triggers to the extracted/generated trigger list, as it leads to a drop in model performance. This can be attributed to the presence of duplicate information as the trigger generation/extraction already accounts for the $k$ triggers. For passage generation/Weak Sup, we append the $k$ datapoints to the synthetically generated data to provide signals from the gold data.
@@ C.3 Downstream Model Training
We choose DEGREE Hsu et al. (2022) as our downstream model for evaluation, a generation-based prompting model that utilizes natural language templates. We implemented the DEGREE model under TextEE framework Huang et al. (2024). Table 11 presents the primary hyperparameters for this model.
@@ C.4 LLM Fine-tuning
We discuss domain-adapted passage generation through LLM fine-tuning in § 6.4. Specifically, we conduct a low-rank finetuning (LoRA) Hu et al. (2021) to reduce computational overhead to fine-tune Llama3-8B-Instruct. We implement LoRA using the peft and trl packages Mangrulkar et al. (2022); von Werra et al. (2020). We choose the task of causal language modeling (i.e. continual pre-training) to perform domain adaptation on unlabeled in-domain sentences. We utilize cross-entropy loss on the dev split of the unlabeled data to select the best model. We provide additional details about the hyperparameters for this fine-tuning for each dataset in Table 12 below.
@@ Appendix D Additional analyses
Figure 5: Model performance for FIG as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
@@ D.1 STAR with domain-specific prompt
A simple way to infuse domain specific information in inverse generation pipelines like STAR would be to add domain-related information in the prompts to the LLM. We experiment with two such methods: (1) domain-mention, where we provide the target domain information in the prompt and ask the model to generate accordingly, and (2) domain-reference, where we use some examples from the unlabeled data in the prompt as reference sentences to better guide the passage generation. We provide results for these explorations using the Llama3-8B-Insturct model in Table 13. As observed, the results are generally poor, with an average drop of 0.1-0.6% F1 for domain-mention and 3.1-3.8% F1 for domain-reference. This is majorly because LLMs over-compensate, producing longer and more stereotypical information in their generations which hurts the naturalness of the sentence and causes further domain drift. Furthermore, the LLM makes more errors in mentioning the event as a part of its reasoning, which is utilized to make the generation in the domain style. We provide some qualitative examples for such generations in Table 14. In some ways, it also puts into light and amplifies the gains obtained by doing target domain SFT for FIG as discussed in § 6.4.
Figure 6: Prompt for stage 1 of forward generation for target trigger extraction.
Figure 7: Prompt for stage 2 of forward generation for target trigger extraction.
Figure 8: Prompt for inverse generation for passage generation.
Figure 9: Prompt for forward generation for data refinement.
@@ D.2 Impact of different number of training samples
We conduct a small analysis to study the impact of changing the number of generated samples on the downstream model performance for FIG. We present the results for Llama3-8B-Instruct in Figure 5. As observed, the performance continues to increase as we increase the data from $N=10$ to $N=100$ datapoints per event type. This promises that inverse generation will provide continued improvements by having larger control over data distribution.
@@ D.3 Human Evaluation Details
We majorly evaluate on three dimensions: (1) Sentence naturalness (SN): this metric judges whether the sentence seems grammatical, natural, and fits the domain of the target data. (2) Event Relevance (ER): this metric is computed only for inverse/hybrid generation methods. This evaluation judges whether the sampled event and trigger are appropriately used to generate a sensible alignment with the target domain. Furthermore, it is verified if the right event definition is used. (3) Annotation Quality (AQ): this metric judges if the right trigger is used for each event mentioned in the synthetic output. If there are any missing events, then this score is penalized. For each metric, a score is given on a Likert scale Likert (1932) from 1 (worst) to 5 (best). We also provide event definitions for each event in each dataset as a reference for better judgment. We illustrate the annotation interface in Figure 10 and provide some sample examples in Table 15.
@@
FIG shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
@@
FIG killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
@@
FIG bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
@@
Table 16: Comparison of generated triggers and sentences from STAR and FIG methods
D.4 Additional Qualitative Examples
In § 6.5, we discussed how FIG improves domain drift qualitatively relative to STARand provided some examples. Here, we provide more examples to further support that study in Table 16. This table further demonstrates how STAR can have a domain drift without a lack of domain-specific cues, while FIG is better here.

### v2

@@ Appendix B Data Statistics
Multilingual Test Data
ACE - Arabic 313 198 24.6
ACE - Chinese 486 211 44.2
Unlabeled Multlingual Train Data
ACE - Arabic 3,218 - 26.1
ACE - Chinese 6,301 - 45.5
Table 10: Data Statistics for the various test and unlabeled datasets used in our work. # = Number of.
@@ Appendix C Implementation Details
Here, we provide detailed implementation details for each component and the models used in our work. We run most of our experiments on NVIDIA RTX A6000/A100 machines with support for 8 GPUs, while for GPT3.5, we make API calls through OpenAI using Curator Marten et al. (2025).
@@ C.1 LLM-based Generation
We provide details on the various hyperparameters for using LLMs in all the components of STAR and SNaRe. For Llama3-8B-Instruct and Llama3-70B-Instruct, we present the hyperparameters in Table 11; while Table 12 presents the hyperparameters for GPT3.5.
@@ C.2 Few-shot Implementation Details
For the few-shot setting, we can access additional $k$ datapoints per event type to aid better performance. For LLM-based prompting, we simply add these examples in the prompt as in-context examples to help the model do better reasoning/generation. For STARand SNaRe, we do not add the $k$ triggers to the trigger list, as it led to a drop in model performance. This can be attributed to the presence of duplicate information, as the trigger generation/extraction already accounts for these gold triggers. We also append the $k$ datapoints to the synthetically generated data to provide signals from the gold data.
@@ C.3 Downstream Model Training
We choose DEGREE Hsu et al. (2022) as our downstream model for evaluation, a generation-based prompting model that utilizes natural language templates. We implemented the DEGREE model under the TextEE framework Huang et al. (2024). Table 13 presents the primary hyperparameters for this model.
@@ C.4 LLM Fine-tuning
We discuss domain-adapted passage generation through LLM fine-tuning in § 7.5. Specifically, we conduct a low-rank finetuning (LoRA) Hu et al. (2021) to reduce computational overhead to fine-tune Llama3-8B-Instruct. We implement LoRA using the peft and trl packages Mangrulkar et al. (2022); von Werra et al. (2020). We choose the task of causal language modeling (i.e., continual pre-training) to perform domain adaptation on unlabeled in-domain sentences. We utilize cross-entropy loss on the dev split of the unlabeled data to select the best model. We provide additional details about the hyperparameters for this fine-tuning for each dataset in Table 14 below.
@@ Appendix D Additional analyses
Figure 5: Model performance for SNaRe as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
@@ D.1 STAR with domain-specific prompt
A simple way to infuse domain-specific information in past works like STAR would be to add domain-related information in the prompts to the LLM. We experiment with two such methods: (1) domain-mention, where we provide the target domain information in the prompt and ask the model to generate accordingly, and (2) domain-reference, where we use some examples from the unlabeled data in the prompt as reference sentences to better guide the passage generation. We provide results for these explorations using the Llama3-8B-Insturct model in Table 15. As observed, the results are generally poor, with an average drop of 0.1-0.6% F1 for domain-mention and 3.1-3.8% F1 for domain-reference. This is mainly because LLMs over-compensate, producing longer and more stereotypical information in their generations, which hurts the naturalness of the sentence and causes further domain drift. Furthermore, the LLM makes more errors in mentioning the event as a part of its reasoning, which is utilized to make the generation in the domain style. We provide some qualitative examples for such generations in Table 16. In some ways, it also puts into light and amplifies the gains obtained by doing target domain SFT for SNaRe as discussed in § 7.5.
Figure 6: Prompt for stage 1 of Scout.
Figure 7: Prompt for stage 2 of Scout.
Figure 8: Prompt for Narrator.
Figure 9: Prompt for Refiner.
@@ D.2 Impact of different number of training samples
We perform a small analysis to study the impact of changing the number of generated samples on the downstream model performance for SNaRe. We present the results for Llama3-8B-Instruct in Figure 5. As observed, performance continues to increase as we increase the data from $N=10$ to $N=100$ datapoints per event type. This promises that data generation will provide continued improvements by practicing greater and better control over the data distribution.
@@ D.3 Human Evaluation Details
We majorly evaluate on three dimensions: (1) Sentence naturalness (SN): This metric judges whether the sentence seems grammatical, natural, and fits the domain of the target data. (2) Event Relevance (ER): This metric is computed only for generation methods that generate sentences from labels. This evaluation judges whether the sampled event and trigger are appropriately used to generate a sensible alignment with the target domain. Furthermore, it is verified if the right event definition is used. (3) Annotation Quality (AQ): This metric judges if the right trigger is used for each event mentioned in the synthetic output. If there are any missing events, then this score is penalized. For each metric, a score is given on a Likert scale Likert (1932) from 1 (worst) to 5 (best). We also provide event definitions for each event in each dataset as a reference for better judgment. We illustrate the annotation interface in Figure 10 and provide some sample examples in Table 17.
Method # Unlabeled Data Eve-I Tri-C
Inference - 46.9 41.3
STAR - 50.0 42.3
Weak Sup 100% train 53.2 48.0
Weak Sup 20% train 51.4 46.1
Weak Sup 5% train 50.1 44.9
SNaRe 100% train 58.1 53.8
SNaRe 20% train 55.9 52.1
SNaRe 5% train 55.1 51.2
Table 18: Ablating the amount of unlabeled data utilized for the ACE dataset by the different data generation methods using Llama3-70B-Instruct and its impact on downstream model performance.
D.4 Analysis in low-resource unlabeled data settings
In rare cases of hyper-specialized domains or domains with privacy concerns, there is a possibility of having super-low availability of unlabeled data. We believe that our current experimentation already simulates this setting, where we use only 773 documents for GENIA and 1.6k tweets for SPEED. Secondly, by providing strong efficacy of external data relative to in-distribution data, we demonstrate how even mildly related unlabeled data can also work well for SNaRe. Finally, we provide a deeper study of varying the amount of unlabeled data for the ACE dataset for Llama3-70B-Instruct as an ablative study to further substantiate the efficacy of our method in Table 18. Lowering the number of unlabeled data reduces the model performance relative to using higher number of data samples. However, even with 5% unlabeled training data ($\sim$850 samples), SNaRe still outperforms the other baselines, showing the efficacy of our method.
@@ D.4 Analysis in low-resource unlabeled data settings
SNaRe shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
@@
SNaRe killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
@@
SNaRe bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
@@
Table 19: Comparison of generated triggers and sentences from STAR and SNaRe methods
D.5 Additional Qualitative Examples
In § 7.6, we discussed how SNaRe improves domain drift qualitatively relative to STARand provided some examples. Here, we provide more examples to further support that study in Table 19. This table further demonstrates how STAR can have a domain drift owing to a lack of domain-specific cues, while SNaRe is better grounded in the target domain.